5.22×10 - 4 mole of a gas containing H2, O2, and N2 exerted a pressure of 67.4 mm in a certain standard volume. The gas was passed over a hot platinum filament which combined H2 and O2 into H2O which was frozen out. When the gas was returned to the same volume, the pressure was 14.3 mm. Extra oxygen was added to increase the pressure to 44.3 mm. The combustion was repeated, after which the pressure read 32.9 mm. What was the mole fraction of H2, O2, and N2 in the gas sample?

 


5.22×10-4 mole of a gas containing H2, O2, and N2 exerted a pressure of 67.4 mm in a certain standard volume. The gas was passed over a hot platinum filament which combined H2 and O2 into H2O which was frozen out. When the gas was returned to the same volume, the pressure was 14.3 mm. Extra oxygen was added to increase the pressure to 44.3 mm. The combustion was repeated, after which the pressure read 32.9 mm. What was the mole fraction of H2, O2, and N2 in the gas sample?

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Nikit,

let H2 and O2 be x and y moles

let given moles be n

n....67.4 mm

n-y-2y.....14.3

1-3y/n = 14.3/67.4

let z moles of O2 is added then

z......30 mm

n-3y-(x-2y)+z-(x/2-y)......32.9

n-3x/2 + z..........32.9

n-3x/2......2.9

1-3x/2n=2.9/67.4

therefore

 3y/n=1-14.3/67.4

 y/n =0.2626

x/n =0.638

molefractions of H2,O2,N2 are 0.638, 0.2626, 0.0994 respectively